Two opposite sides of a square with length x are increased by 3 units each and the other two opposite sides are decreased by 3 units each. If the area of the resulting rectangle is 27 units, what would be the width of the rectangle?
Hint: area of rectangle = length × width

Answer by Theo(13342)

x equals the length of the side of the square.
length of the rectangle is equal to x+3
width of the rectangle is equal to x-3
area of the rectangle is equal to length * width which becomes:
(x-3) * (x+3) = 27
multiply the factors together and you get:
x^2 - 9 = 27
add 9 to both sides of the equation to get:
x^2 = 36
take the square root of both sides of the equqtion to get:
x = plus or minus 6
minus is no good.
x = 6
x+3 = 9
x-3 = 3
length of the rectangle is 9
width of the rectangle is 3.